OpenStudy (anonymous):

Need help setting up equations ! You found a bag of dollar bills and you return it to the police. Later on, an officer tells you that the bag contained 3 kinds of bills an O amount of $1 bills, F amount of $5, and L amount of $50 dollar bills. The bag had 110 bills worth $1620. Moreover, there were 3 times as many $50 bills than $5 bills. Write a system of 3 linear equations. .. I'm pretty sure the first two are O+F+L=110, and O+5F+50L=1620, but I can't figure out the last one.

OpenStudy (anonymous):

3(L)+F=100-O

OpenStudy (anonymous):

Sorry 110-O cause you have 110 bills

OpenStudy (anonymous):

how would I set it up to where I can put it in a matrix though?

OpenStudy (anonymous):

I would think 3L + F + 0 = 110

OpenStudy (anonymous):

:/ it doesn't work cuz when I put all the equations in, it gives me [1 0 0 -267.5 0 1 0 377.5 0 0 1 0] which just doesnt really make sense

OpenStudy (anonymous):

Let me look at it some and I'll get back to you. It's bout time for me to go tonight.